OpenStudy (anonymous):

Candies are packed in four different sized packets containing 200,250,350,800 candies respectively. Find the smallest number of candies needed to make an exact number of packets of each size. Also find the number of packets of each kind that will be made from this number

OpenStudy (anonymous):

So you have to break each number down into it's prime number and then see which ones are similar between the two. Then you have to multiply that number out to get 200, 250, 350 & 800

OpenStudy (noelgreco):

Find the LCM, and then apportion the candies. Hint: the LCM is 28,000.

OpenStudy (anonymous):

Don't you need to find the greatest common denominator and then find how many packets you need to make the total?

OpenStudy (anonymous):

Breaking them down 200 = 5^2 * 2^3 250 = 5^3 * 2 350 = 7 * 5^2 * 2 800 = Then find out what is shared by all of them and multiply those together to get your greatest common factor (sorry) Then divide each total candy bag by your GCF to get the 2nd part of the answer.